OpenStudy (anonymous):

Rewrite the following in Ax + By = C form: a. y=2x - 1 b. y=-9x - 5 c. y=6x - 6

OpenStudy (anonymous):

help

OpenStudy (bobbyleary):

In Ax + By = C: A,B, and C are coefficients. So for a. you must bring the 2x to the left side. -2x + y = -1 The others are done similarly. Post your solutions for b and c and I can double check them for you.

OpenStudy (anonymous):

b. 9x + y = -5 c. -6x + y = -6 right???

OpenStudy (bobbyleary):

Yep. Perfect!

OpenStudy (anonymous):

Thanks